A box contains 6 nails and 10 nuts. Half of the nails and half of the nuts are rusted. If one item is chosen at random, then the probability that it is rusted or a nail is.........

Grade:11

2 Answers

Aman Bansal
11 years ago

Dear Shiva,

P(rusted) = 8/16

P(nail) = 6/16

P(rusted and nail) = 3/16

P(A intersection B) = 8/16 + 6/16 - 3/16;

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. 

http://www.askiitians.com/packages/packages.aspx

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

upender surepally
11 years ago

11/16